Fermi National Laboratory


Accelerator Update

Friday August 22
The day shift began with Operations monitoring store 2934, with the Antiprotons Source (Pbar) stacking, with Recycler experts conducting studies, and with MiniBooNE taking beam.

At 8:15 AM, Linac experts began working on RF station 3 and other problems.

Operations terminated store 2934 at 8:34 AM.

The Collider Detector Facility (CDF) and D-zero (D0) accessed their collision halls around 9 AM. CDF and D0 were out of access by 12 PM.

Linac beam came back at 11:59 AM. Stacking and beam to MiniBooNE resumed.

Linac RF stations 3&4 began suffering from intertank phasing problems again at 1 PM.

At 2:06 PM, operators noted a LCW leak alarm for the Main Injector (MI) RF. They contacted a water expert.

Tevatron (TeV) experts began a study period at 2:21 PM.

Switchyard experts began a study period at 4:15 PM.

LRF3 tripped off at 5:42 PM. It reset okay.

Operations put the TeV into shot setup and established store 2937 at 8:38 PM. Its initial luminosity was 4.1E31. Pbar resumed stacking.

Saturday August 23
The midnight shift began with Operations monitoring store 2937, with Pbar stacking, and with MiniBooNE taking beam.

A Switchyard expert began a study period at 9:09 AM.

A MI expert began a study period at 9:56 AM.

A Booster RF station tripped off at 10:02 AM. It reset okay. Then stations 5&6 tripped off on a vacuum burst at 10:11 AM. Beam was back two minutes later.

BRF18 tripped off at 10:39 AM. It reset okay.

At 12:10 PM, MI began suffering from beam problems. Operators worked at tuning the accelerator, but the problems still kept MiniBooNE off. At 1:48 PM, Operations discovered a Booster damper that had tripped off. It reset okay. This damper tripped off three more times between 3 and 3:30 PM. Later this was discovered to be a temperature problem in the equipment room. Operators installed a fan.

At 3 PM, MI RF stations 5, 13, & 14 tripped off. They reset okay, but they tripped off three more times over the next hour and a half.

Operations terminated store 2937 at 4:02 PM.

The Switchyard expert completed his study at 4:15 PM.

CDF and D0 accessed their collision halls at 4:25 PM.

MI studies finished at 5:02 PM.

D0 came out of access at 5:04 and CDF came out at 5:26 PM.

Operations put the TeV into shot setup at 6:10 PM and at 8:07 PM established store 2939 with an initial luminosity of 4.2E31. Pbar resumed stacking.

A Linac RF station (LRF2) tripped off at 11:54 PM. It reset okay.

Sunday August 24
The midnight shift began with Operations monitoring store 2939, with Pbar stacking, and with MiniBooNE taking beam.

A Booster RF station (BRF4) tripped off at 6:10 AM. It reset okay,

A Linac RF station (LRF4) tripped off at 7:15 AM and then again three minutes later. It reset okay both times. LRF1 also tripped off during this time.

BRF4 tripped off at 7:30 AM. It reset okay.

BRF5 & 6 tripped off due to high vacuum pressure. They reset okay.

LRF3 tripped off at 8:54 AM. It reset okay.

LRF4 tripped off at 9:55 AM. It reset okay.

A NuMI expert began a study at 11:04 AM.

MIRF3, 10, & 13 tripped off at 11:23 AM. They rest okay.

Operations terminated store 2939 at 2:13 PM.

MI studies ended at 2:30 PM.

Operations put the TeV into shot setup at 3:04 PM. At 4:47 PM, Operations established store 2941 with an initial luminosity of 4.0E31. Pbar resumed stacking.

Operations tuned the Booster beam from 6:30 to 7:30 PM. Operations started tuning the MI beam at 8:46 PM.

Monday August 25
The midnight shift began with Operations monitoring store 2941, with Pbar stacking, and with MiniBooNE taking beam.

At 2:24 AM, a power glitch caused the TeV to quench at sector B0. Many devices tripped off all over the complex, including wet engines and cryo systems. Operations contacted the Duty Electrician and he said there was problem with feeder 47.

Operations established beam to Pbar at 3:15 AM, and beam to MiniBooNE at 3:41 AM.

Today's Plan
The plan is to establish another store as soon as possible this morning. There will be a TeV and Pbar study period on Tuesday. Thursday will be a shutdown day to perform normal maintenance and to prepare for the big shut down next month.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 8/25/2003   email Fermilab